OpenStudy (anonymous):

I really need help fan, medal, and testimony Write an equation of a line that passes through the point (2, 3) and is parallel to the line y = 3 over 2x + 5.

OpenStudy (anonymous):

@satellite73 @ganeshie8 @sleepyjess

OpenStudy (anonymous):

Ok, are you familiar with the line equation \(y = mx + b\)?

OpenStudy (anonymous):

slope?

OpenStudy (anonymous):

\(m\) is the slope of the line The other line is \(y = \frac{3}{2}x + 5\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so since the lines are parallel we know they must have the same slope. The slope of the second line is \(\frac{3}{2}\) so our line is now: $$ y = mx + b \implies y = \frac{3}{2}x + b$$ Now we are told it has to have the point (2,3). That means that when x is 2 y has to be 3. Let's plug it in: $$ x = 2 \implies y = 3\\ y = \frac{3}{2}x + b \implies 3= \frac{3}{2} \cdot 2 + b $$Can you solve for \(b\)?